OpenStudy (anonymous):

solve equation write mathematical model representing the monthly cost, C, there are two options, let x=hours per month of court rental time #1, option35$ monthly fee plus $4hour for court rental#2 option, no monthly fee but $6.50hour/hour for court rental

OpenStudy (anonymous):

solve equation write mathematical model representing the monthly cost, C, there are two options, let x=hours per month of court rental time #1, option35$ monthly fee plus $4hour for court rental#2 option, no monthly fee but $6.50hour/hour for court rental how many rentals is option1 less than option2

OpenStudy (anonymous):

my full question in attachment so you can get better understanding

jimthompson5910 (jim_thompson5910):

Were you able to find any of the equations?

OpenStudy (anonymous):

no can u look at the attachment

OpenStudy (anonymous):

the option1 and2 is whati need help on

jimthompson5910 (jim_thompson5910):

ok if x = hours per month of court rental time and C = total cost per month then the first equation is C = 4x + 35 Why? because if you rent the court for x hours, then you pay $4 an hour, and you pay 4x total...but you also pay the fixed monthly fee on top of that which is $35 So you pay a total of 4x+35 dollars which is why the first equation is C = 4x + 35

jimthompson5910 (jim_thompson5910):

the second option is no monthly fee, so the second equation is just C = 6.5x

jimthompson5910 (jim_thompson5910):

So we have C = 4x + 35 C = 6.5x Set them equal to each other to get 6.5x = 4x + 35 Now solve for x

jimthompson5910 (jim_thompson5910):

does that make sense?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (anonymous):

now do i have to distribute the like terms

jimthompson5910 (jim_thompson5910):

6.5x = 4x + 35 6.5x - 4x = 35 ... Subtract 4x from both sides. Keep going

OpenStudy (anonymous):

31[6.5

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

6.5x = 4x + 35 6.5x - 4x = 35 2.5x = 35 x = 35/2.5 x = 14

jimthompson5910 (jim_thompson5910):

This means that the two costs will be the same when you rent the court for 14 hours total in a month

jimthompson5910 (jim_thompson5910):

Anything past 14 hours will make the second plan (the one without the monthly fee) more expensive

jimthompson5910 (jim_thompson5910):

Anything below 14 hours will make the first plan (with the monthly fee) more expensive

OpenStudy (anonymous):

okay thanku

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now working on another

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

this is what i am working on now and is clueless

jimthompson5910 (jim_thompson5910):

have you gotten started at all?

OpenStudy (anonymous):

yes but still confused

OpenStudy (anonymous):

the base is 3 oflog35=y loga7=-2=a-2=7

jimthompson5910 (jim_thompson5910):

\[\Large \log_{3}(5) = y\] is the same as \[\Large 3^y = 5\] using the rule provided in the pdf

jimthompson5910 (jim_thompson5910):

\[\Large \log_{a}(7) = -2\] is the same as \[\Large a^{-2} = 7\] using the rule provided in the pdf and you're doing that with the rest

OpenStudy (anonymous):

no because it say solve

OpenStudy (anonymous):

for partb

jimthompson5910 (jim_thompson5910):

you would use the rule used in part a to isolate the variable

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so log2x-3

jimthompson5910 (jim_thompson5910):

that would turn into x = 2^(-3) x = 1/(2^3) x = 1/8

OpenStudy (anonymous):

so logx9=1/2=x=1/2(9)

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

or x=1/(9^1/2)=2/1

jimthompson5910 (jim_thompson5910):

x = 9^(1/2) x = sqrt(9) x = ???

OpenStudy (anonymous):

81

jimthompson5910 (jim_thompson5910):

the square root of 9 NOT 9 squared

OpenStudy (anonymous):

ok,3

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

for 2 the graph it would be 10 left1 right

jimthompson5910 (jim_thompson5910):

\[\Large e^{x}+2\] shifts the entire graph up 2 units

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hx=-ex=0

jimthompson5910 (jim_thompson5910):

sticking a negative on e^x flips the graph over the x-axis

OpenStudy (anonymous):

so it would be-e^x

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (anonymous):

for3a do i just add

jimthompson5910 (jim_thompson5910):

you plug in x = 5 and evaluate

OpenStudy (anonymous):

not sure done right i got112

jimthompson5910 (jim_thompson5910):

way too high

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-3=12-20+5=-3

jimthompson5910 (jim_thompson5910):

do you have a calculator?

OpenStudy (anonymous):

not working

jimthompson5910 (jim_thompson5910):

use google as a calculator then http://www.google.com/search?hl=&q=12%2B20*ln%285%29&sourceid=navclient-ff&rlz=1B3GGLL_enUS420US420&ie=UTF-8

OpenStudy (anonymous):

ok so i have to plug inp=12+20/5

OpenStudy (anonymous):

ok thanks so it would be the same equation but would have to multiply6

OpenStudy (anonymous):

44. 1887582487 but have to round to six decimals so it would be 50

jimthompson5910 (jim_thompson5910):

it would be 44.188758 after rounding

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if ,i were to add 2002+2005=4007, but how would i begin to solve for 2005

jimthompson5910 (jim_thompson5910):

why are you adding 2002 to 2005 ?

OpenStudy (anonymous):

i thought i have to

jimthompson5910 (jim_thompson5910):

no you don't, I'm not exactly sure what you're doing

OpenStudy (anonymous):

me either but trying to solve for2005

jimthompson5910 (jim_thompson5910):

you're not solving for 2005, you're plugging in x =5 and evaluating

OpenStudy (anonymous):

12+20xin5=44.18 round two decimals

jimthompson5910 (jim_thompson5910):

it's Ln not in, and you should have it rounded to 6 decimal places

OpenStudy (anonymous):

44.188758

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so if i have to round it to more decimals it would be44.288758

jimthompson5910 (jim_thompson5910):

just leave it at 44.188758 and that's fine

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

b, would be 90%=0.9

jimthompson5910 (jim_thompson5910):

solve 12+20*ln(x) = 90 for x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can i use 10 or just1 in the model?

jimthompson5910 (jim_thompson5910):

what do you mean 10?

OpenStudy (anonymous):

90 then idid 12+20*Ln(90)=101.996193407

OpenStudy (anonymous):

okay, i just use 1

jimthompson5910 (jim_thompson5910):

no you solve 12+20*ln(x) = 90 for x

jimthompson5910 (jim_thompson5910):

you don't plug in 90 for x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

14

OpenStudy (anonymous):

12+20*Ln(-8)=90

jimthompson5910 (jim_thompson5910):

12+20*ln(x) = 90 20*ln(x) = 90-12 20*ln(x) = 78 ln(x) = 78/20 ln(x) = 3.9 x = ???

OpenStudy (anonymous):

78

OpenStudy (anonymous):

78/20=3.9=x

OpenStudy (anonymous):

12

OpenStudy (anonymous):

because x=90

OpenStudy (anonymous):

3900 because i have to round 2 decimals